# The Artificial Benchmark for Community Detection with Outliers and Overlapping Communities (ABCD+o2)

### v1

@@ 5 Conclusion
We are also interested in results for the ABCD+o2 model that generalize results of the ABCD and ABCD+o models. In [19] the modularity was studied and it was found that the maximum modularity came from the ground truth communities until a certain level of noise, after which a higher modularity could be attained. A similar behaviour should be seen with ABCD+o2 and its overlap parameter $\eta$, although we would need to select an appropriate generalization of modularity for overlapping sets, such as the generalization proposed in [12]. Additionally, in [5] it was shown that ABCD graphs exhibit self-similar behaviour, namely, the degree distributions of communities are asymptotically the same as the degree distribution of the whole graph (up to an appropriate normalization). We suspect that this self-similar property persists in ABCD+o2.

### v2

@@ 5 Conclusion
We are also interested in results for the ABCD+o2 model that generalize results of the ABCD and ABCD+o models. In [kaminski2022modularity] the modularity was studied and it was found that the maximum modularity came from the ground truth communities until a certain level of noise, after which a higher modularity could be attained. One could test for a similar result in ABCD+o2 by comparing the maximum modularity to both the level of noise $\xi$ and the overlap parameter $\eta$. A first step to this future research would be selecting an appropriate generalization of modularity for overlapping sets, such as the generalization proposed in [devi2016overlapping]. Additionally, in [self-similarityABCD] it was shown that ABCD graphs exhibit self-similar behaviour, namely, the degree distributions of communities are asymptotically the same as the degree distribution of the whole graph (up to an appropriate normalization). Another open problem is checking if this self-similar property persists in ABCD+o2.
@@ Appendix A Density of Intersection Figures
Figures 1112 and 13, together with the previously shown Figure 5, present the full experiment comparing community and intersection densities for various $\rho$.
@@ Appendix B Internal Edge Fraction Figures
Figures 1415 and 16, together with the previously shown Figure 7, present the full experiment showing the range of the top five internal edge fractions, sorted in descending order, on the real networks (left) and the corresponding ABCD+o2 graphs (right), binned by the number of communities the node $v$ is in.